OpenStudy (anonymous):

Mike is saving money for his first car. He already saved $300 from cutting grass last summer. He has an after school job and saves $250 from his paychecks every month. This relationship is represented by the equation y = 250x + 300, where x is the number of months and y is the cost of the car. 1. Identify the slope (m) in this relationship. 2. Explain what the slope (m) means in this situation. 3. Identify the y-intercept (b) in this relationship. 4. Explain what the y-intercept (b) means in this situation. 5. Use the equation to determine the number of months Mike will nee

OpenStudy (anonymous):

Slope: 250

OpenStudy (anonymous):

For every month he works, he makes $250

OpenStudy (anonymous):

y-intercept: 300

OpenStudy (anonymous):

meaning: the initial amount of money he had starting out was $300.

OpenStudy (anonymous):

5. is cut off.

OpenStudy (anonymous):

5. Use the equation to determine the number of months Mike will need to save for a car that costs $4000. 6. Use the equation to determine the maximum amount of money Mike can afford to spend on a car if he saves for 12 months. 7. Explain how the amount of time Mike will take to save for his $4000 car will change if he increased the amount of money he saved each month. 8. Explain how the amount of time Mike will take to save for his $4000 car will change if he decreased the amount of money he saved each month.

OpenStudy (anonymous):

5. 4000 = 250x+300. Solve for x. Your answer will be in months. 6. y = 250(12) +300

OpenStudy (anonymous):

7. Lol he would be able to afford the car faster! 8. It would take longer to save up for the car. (too bad 7 and 8 aren't more specific).

OpenStudy (anonymous):

Thanks For Helping me I was Sooo Confused. (;

OpenStudy (anonymous):

Do you have a good resource as far as your textbook goes?

OpenStudy (anonymous):

Nooopeee.

OpenStudy (anonymous):

Okay lemme see what I can do about that.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

when I took algebra 7 years ago, my book was a total piece of pellet.

OpenStudy (anonymous):

Paul's notes have always helped me for calculus and maybe the algebra ones will help for you: http://tutorial.math.lamar.edu/Classes/Alg/Graphing.aspx

OpenStudy (anonymous):

hahaha they censor on here!

OpenStudy (anonymous):

Haha(: Thanks

OpenStudy (anonymous):

So for this problem, your output (y) is in dollars and your input (x) is in months. So if they say if the amount is however many dollars, you know that you have to plug that in to y. If ask something about months, then you plug that into x.

OpenStudy (anonymous):

When i solve for x on this problem 4000 = 250x+300. do i do 250+300 ?

OpenStudy (anonymous):

4000 - 300 = 250x 3700 = 250x (3700/250) = (250x/250) 3700/250 = x

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

x = 14.8 months. But we're going to round up. So x= about 15 months

OpenStudy (anonymous):

15 months

OpenStudy (anonymous):

Does input and output stuff make sense to you?

OpenStudy (anonymous):

yeah i get it now.

OpenStudy (anonymous):

Yeah Our books Now Aree Stupid too